# PoIFusion: Multi-Modal 3D Object Detection via Fusion at Points of Interest

### v1

@@ 3.3 Feature Sampling
{$m^{i}&=\frac{x^{i}-X_{\text{min}}}{(X_{\text{max}}-X_{\text{min}})\times d},\\ n^{j}&=\frac{y^{i}-Y_{\text{min}}}{(Y_{\text{max}}-Y_{\text{min}})\times d},\end{aligned}\right.$ (2)
@@
Typically, there are multiple surrounding-view cameras on the autonomous vehicle, resulting in multi-view images. A 3D point can be projected onto one or two views. If the PoI is projected to only one view, this view will be leveraged to perform image feature sampling. Otherwise, we follow [5] to randomly choose one view for image feature sampling. Given a 3D PoI $P^{i}$ in the LiDAR coordinate system, the projection onto the 2D image plane can be computed as follows:
@@
In our camera encoder, we exploit an image backbone network followed by an FPN [35] to produce feature maps at P2, P3, P4, and P5 (1/4, 1/8, 1/16, and 1/32 downsampling, respectively). For a given projected point $P_{\text{image}}^{i}$, a set of image feature $\{\bm{g}^{i}_{j}\}_{j=2}^{5}$ is sampled from the multi-scale feature maps via bilinear interpolation. The sampled multi-scale image features are then aggregated as:
@@
where $\bm{f}_{I}^{i}$ represents the aggregated image feature and $\{w^{i}_{j}\}_{j=2}^{5}$ are scale weight coefficients, which are predicted by a linear layer according to the query feature $\bm{v}$.
Figure 3: An illustration of our dynamic fusion block.
@@ 3.4 Dynamic Multi-Modal Feature Fusion
The detail of our dynamic fusion block is depicted in Figure 3. After PoI generation and feature sampling, each object query $\bm{Q}$ is represented by a set of PoIs $\{P^{i}\}$, and each PoI is attached with a sampled multi-modal feature pair $\{\bm{f}_{P}^{i},\bm{f}_{I}^{i}\}$. Firstly, we exploit conventional linear layers to produce dynamic fusion parameters $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ according to the query feature $\bm{v}$. After that, the concatenated multi-modal feature pair of each PoI is individually integrated through the dynamic linear layers parameterized by $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$. Note that the distinction between the dynamic linear layer and the conventional linear layer is that the parameters of the dynamic one are produced on the fly. Subsequently, the features of PoIs derived from the same object query are concatenated together and fed into an additional linear layer to perform PoI feature aggregation. Finally, the aggregated feature is added back to the query feature. The dynamic linear layers and the conventional linear layer leveraged for PoI feature aggregation are followed by a Layer Normalization operation [1] and a ReLU activation layer.
@@ 3.5 Prediction Head and Training Objective
Prediction Head.
Our prediction head consists of a classification head and a sibling box regression head. In the classification head, we predict the binary classification score for each category. In the regression head, we follow the iterative refinement scheme [55, 84, 37] to predict the center delta over the predicted box center from the previous iteration. The other parts of the predicted 3D bounding box, including the box dimension, the heading direction, and the velocity, are independently predicted at each time.
Training Objective.
We follow the practice of set prediction [6] for target assignment. Specifically, we exploit bipartite matching to perform one-to-one assignments between the predicted 3D bounding boxes and the ground-truth boxes. In corresponding to the prediction head, our training objective also includes two parts: a focal loss [36] for the classification head and an L1 loss for the regression head. The overall training objective is computed as:
@@
where $\bm{y}_{pred}$ is the classification score, $\bm{b}_{pred}$ is the predicted 3D box, $\alpha$ and $\beta$ are set as 2.0 and 0.25 following [69, 40] to balance the losses. The training objective is applied to the predicted output of each recurrent iteration.
@@ 4 Experiments
Method Present at Modality Camera Backbone LiDAR Backbone val test
NDS $\uparrow$ mAP $\uparrow$ NDS $\uparrow$ mAP $\uparrow$
DETR3D [63] CORL’21 C ResNet-50 - 43.4 34.9 47.9 41.2
BEVFormer [33] ECCV’22 C ResNet-50 - 51.7 41.6 56.9 48.1
CenterPoint [76] CVPR’21 L - VoxelNet 66.8 59.6 67.3 60.3
TransFusion-L [2] CVPR’22 L - VoxelNet 70.1 65.1 70.2 65.5
${}^{\star}$  [77] NeurIPS’21 C+L ResNet-50 VoxelNet 70.8 67.1 70.5 66.4
PointAugmenting [58] CVPR’21 C+L ResNet-50 VoxelNet - - 71.0 66.8
FUTR3D [8] CVPRW’22 C+L ResNet-101 VoxelNet 68.3 64.5 - -
TransFusion [2] CVPR’22 C+L ResNet-50 VoxelNet 71.3 67.5 71.6 68.9
UVTR [29] NeurIPS’22 C+L ResNet-101 VoxelNet 70.2 65.4 71.1 67.1
BEVFusion (PKU) [34] NeurIPS’22 C+L Dual-Swin-T VoxelNet 72.1 69.6 73.3 71.3
DeepInteraction [75] NeurIPS’22 C+L ResNet-50 VoxelNet 72.6 69.9 73.4 70.8
BEVFusion (MIT) [42] ICRA’23 C+L Swin-T VoxelNet 71.4 68.5 72.9 70.2
${}^{\star}$ [26] CVPR’23 C+L ResNet-50 VoxelNet 72.0 69.3 74.0 71.5
${}^{\star}$ [31] arXiv’23 C+L ResNet-50 Sparse-ResUNet 72.7 70.4 74.0 70.6
UniTR [60] ICCV’23 C+L Multi-Modal DSVT Multi-Modal DSVT 73.1 70.0 74.1 70.5
SparseFusion [67] ICCV’23 C+L ResNet-50 VoxelNet 72.8 70.4 73.8 72.0
${}^{\triangle}$ [69] ICCV’23 C+L VoVNet-99 VoxelNet 72.9 70.3 74.1 72.0
ObjectFusion [5] ICCV’23 C+L Swin-T VoxelNet 72.3 69.8 73.3 71.0
PoIFusion (ours) - C+L ResNet-50 VoxelNet 73.2 71.2 74.3 72.6
PoIFusion (ours) - C+L Swin-T VoxelNet 73.6 71.7 74.9 73.4
Table 1: Comparison of 3D object detection performance on nuScenes dataset. “C” means camera data (RGB images), “L” means LiDAR data (point clouds), $\star$” indicates the image input depends on predicted 2D instance masks, and $\triangle$” indicates the image resolution is enlarged to $1600\times 640$. We highlight the top-2 entries with bold font. We do not apply any test-time augmentation or model ensemble.
@@ 4.1 Experimental Setup
Dataset and Metric.
We evaluate our method on the nuScenes dataset [4], which includes 700, 150, and 150 driving scenes for training, validation, and testing. The vehicle for data collection is equipped with a 32-beam LiDAR sensor and 6 surrounding-view RGB cameras, thus providing both point clouds and multi-view images. Besides, this dataset annotates more than 1.4 million 3D bounding boxes, across 10 common categories on the street. Both the multi-sensor data and annotation facilitate the exploration of multi-modal 3D object detection. We follow the official policy to mainly evaluate our method in terms of mean average precision (mAP) and nuScenes detection score (NDS). The mAP is averaged over distance thresholds $0.5m$, $1m$, $2m$, and $4m$ on the BEV across $10$ classes. NDS is a weighted average of mAP and other true-positive metrics including mATE, mASE, mAOE, mAVE, and mAAE.
Network Configuration.
The camera encoder is composed of an image backbone network (i.e., ResNet [23] or Swin-Transformer [41]) and an FPN [35] to enrich the multi-scale information. The LiDAR encoder consists of a 3D sparse VoxelNet [80, 70] and a BEV backbone network. We follow the common practice of setting the image resolution as $800\times 448$ and setting the voxel size as ($0.075m$, $0.075m$, $0.2m$). In our multi-modal object decoder, there are $900$ object queries. The feature dimension of query embedding is set as $256$. The channel of image and point cloud features is also transformed to $256$ before being fed into the multi-modal decoder. We equally divide the channels of feature maps of each modality into 4 groups, and generate PoIs for each group. This grouping operation improves the capacity of the network [66] and reduces the computation costs of the dynamic fusion block. Moreover, the first dynamic fusion layer halves the channel of the query embedding, and the second dynamic fusion layer restores the feature dimension. The multi-modal object decoder is iteratively applied 6 times, with shared parameters.
Training and Inference.
We follow the same training pipeline of previous methods [42, 5, 67, 34, 75], and use the same pretrained models. Specifically, the camera encoder is initialized with the model weights of a Mask R-CNN [24] model pre-trained on nuImage [4] dataset. The LiDAR encoder is initialized with the model weights of pretrained TransFusion-L [2]. Our fusion framework is trained for 6 epochs with the AdamW optimizer. The training sample is resampled by CBGS [82] strategy. The initial learning rate is set as 1e-4, adapted with the one-cycle learning rate policy, and the weight decay is set as 0.01. To speed up training convergence, the query denoising technique [39, 69, 38] is utilized. Data augmentation including random flip, random rotation, random translation, random scaling, and random modal masking is adopted. We exploit 8 GPUs for training, with 2 samples on each GPU.
@@
4.2 Comparison with Other Methods
Main Results of Detection.
@@ 4.2 Comparison with Other Methods
In Table 1, we conduct a comprehensive comparison on the nuScenes 3D object detection benchmark. The methods are divided into three groups: camera-based (“C”), LiDAR-based (“L”), and multi-modal (“C+L”). Broadly speaking, the multi-modal methods outperform the methods that leverage only one single modality, validating the benefits of integrating the semantic-intensive RGB image and location-aware LiDAR point clouds.
Method
Modality
AMOTA (%) $\uparrow$
AMOTP (%) $\downarrow$
CenterPoint [76]
L
63.7
60.6
VoxelNeXt [11]
L
70.2
64.0
TransFusion [2]
C+L
71.8
60.3
BEVFusion (MIT) [42]
C+L
72.8
59.4
ObjectFusion [5]
C+L
74.2
54.3
PoIFusion(ours)
C+L
75.1 50.7

### v2

@@ 3.4 Feature Sampling
$\begin{cases}m^{i}&=\dfrac{x^{i}-X_{\text{min}}}{\left(X_{\text{max}}-X_{\text{min}}\right)\times d},\\[8.61108pt] n^{j}&=\dfrac{y^{i}-Y_{\text{min}}}{\left(Y_{\text{max}}-Y_{\text{min}}\right)\times d},\end{cases}\end{aligned}$ (6)
@@
Typically, there are multiple surrounding-view cameras on the autonomous vehicle, resulting in multi-view images. A 3D point can be projected onto one or two views. If the PoI is projected to only one view, this view will be leveraged to perform image feature sampling. Otherwise, we follow Cai et al. (2023) to randomly choose one view for image feature sampling. Given a 3D PoI $P^{i}$ in the LiDAR coordinate system, the projection onto the 2D image plane can be computed as follows:
@@
In our image encoder, we exploit an image backbone network followed by an FPN Lin et al. (2017a) to produce feature maps at P2, P3, P4, and P5 (1/4, 1/8, 1/16, and 1/32 downsampling, respectively). For a given projected point $P_{\text{image}}^{i}$, a set of image feature $\{\bm{g}^{i}_{j}\}_{j=2}^{5}$ is sampled from the multi-scale feature maps via bilinear interpolation. The sampled multi-scale image features are then aggregated as:
@@
where $\bm{f}_{I}^{i}$ represents the aggregated image feature and $\{w^{i}_{j}\}_{j=2}^{5}$ are scale weight coefficients, which are predicted by a linear layer with the query feature $\bm{v}$.
@@ 3.5 Dynamic Multi-Modal Feature Fusion
The detail of our dynamic fusion block is depicted in Figure 3. After PoI generation and feature sampling, each object query $\bm{Q}$ is represented by a set of PoIs $\{P^{i}\}$, and each PoI is attached with a sampled multi-modal feature pair $\{\bm{f}_{P}^{i},\bm{f}_{I}^{i}\}$. Firstly, we exploit conventional linear layers to produce dynamic fusion parameters $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ according to the query feature $\bm{v}$. After that, the concatenated multi-modal feature pair of each PoI is individually integrated through the dynamic linear layers parameterized by $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$. Note that the distinction between the dynamic linear layer and the conventional linear layer is that the parameters of the dynamic one are produced on the fly. Subsequently, the features of PoIs derived from the same object query are concatenated together and fed into an additional linear layer to perform PoI feature aggregation. Finally, the aggregated feature is added back to the query feature. The dynamic linear layers and the conventional linear layer leveraged for PoI feature aggregation are followed by a Layer Normalization operation Ba et al. (2016) and a ReLU activation layer.
Figure 3: An illustration of our dynamic multi-modal feature fusion block, which first fuses image and point cloud feature at each PoI, and then integrates the PoIs of the same object query in canonical order.
@@ 3.6 Prediction Head and Training Objective
Prediction Head. Our prediction head consists of a classification head and a sibling box regression head. In the classification head, we predict the binary classification score for each category. In the regression head, we follow the iterative refinement scheme Teed and Deng (2020); Zhu et al. (2020); Lin et al. (2023) to predict the center delta over the predicted box center from the previous iteration as follows:
$\bm{c}_{t}=\bm{c}_{t-1}+\Delta\bm{c}_{t},$ (9)
where $\bm{c}_{t}$ is the predicted box center predicted in the $t$-th recurrent iteration and $\Delta\bm{c}_{t}$ is the center delta. The other parts of the predicted 3D bounding box, including the box dimension, the heading direction, and the velocity, are independently predicted at each time.
Training Objective. We follow the practice of set prediction Carion et al. (2020) for the target assignment. Specifically, bipartite matching is exploited to obtain one-to-one assignments between predicted results and ground-truth ones. In corresponding to the prediction head, our training objective also includes two parts: a focal loss Lin et al. (2017b) for the classification head and an L1 loss for the regression head. The overall training objective is computed as follows:
@@
where $\bm{y}$ indicates classification logits, $\bm{b}$ means box coordinates, and $\alpha$ and $\beta$ are coefficients that balance these two losses. We set $\alpha$ to 2.0 and set $\beta$ to 0.25 following Yan et al. (2023); Liu et al. (2022).
@@ 4 Experiments
Table 1: Performance comparison with state-of-the-art methods on nuScenes validation/test split. C: camera data (RGB images). L: LiDAR data (point clouds). $\star$: the image input depends on predicted 2D instance masks. $\triangle$: the image resolution of CMT is enlarged to $1600\times 640$. Our approach does not utilize test-time augmentation or employ model ensembling.
Methods Modality Camera Backbone LiDAR Backbone val test
mAP $\uparrow$ NDS $\uparrow$ mAP $\uparrow$ NDS $\uparrow$
DETR3D Wang et al. (2022) C ResNet-50 - 34.9 43.4 41.2 47.9
BEVFormer Li et al. (2022c) C ResNet-50 - 41.6 51.7 48.1 56.9
CenterPoint Yin et al. (2021a) L - VoxelNet 59.6 66.8 60.3 67.3
TransFusion-L Bai et al. (2022) L - VoxelNet 65.1 70.1 65.5 70.2
MVP  Yin et al. (2021b) C+L ResNet-50 VoxelNet 67.1 70.8 66.4 70.5
PointAugmenting Wang et al. (2021a) C+L ResNet-50 VoxelNet - - 66.8 71.0
FUTR3D Chen et al. (2022a) C+L ResNet-101 VoxelNet 64.5 68.3 - -
TransFusion Bai et al. (2022) C+L ResNet-50 VoxelNet 67.5 71.3 68.9 71.6
UVTR Li et al. (2022a) C+L ResNet-101 VoxelNet 65.4 70.2 67.1 71.1
BEVFusion (PKU) Liang et al. (2022) C+L Dual-Swin-T VoxelNet 69.6 72.1 71.3 73.3
DeepInteraction Yang et al. (2022c) C+L ResNet-50 VoxelNet 69.9 72.6 70.8 73.4
BEVFusion (MIT) Liu et al. (2023b) C+L Swin-T VoxelNet 68.5 71.4 70.2 72.9
MSMDFusion Jiao et al. (2023) C+L ResNet-50 VoxelNet 69.3 72.0 71.5 74.0
UniTR Wang et al. (2023b) C+L Multi-Modal DSVT Multi-Modal DSVT 70.0 73.1 70.5 74.1
SparseFusion Xie et al. (2023) C+L ResNet-50 VoxelNet 70.4 72.8 72.0 73.8
CMT Yan et al. (2023) C+L VoVNet-99 VoxelNet 70.3 72.9 72.0 74.1
ObjectFusion Cai et al. (2023) C+L Swin-T VoxelNet 69.8 72.3 71.0 73.3
FSF Li et al. (2024) C+L ResNet-50 Sparse-ResUNet 70.4 72.7 70.6 74.0
PoIFusion (ours) C+L ResNet-50 VoxelNet 71.2 73.2 72.6 74.3
PoIFusion (ours) C+L Swin-T VoxelNet 71.7 73.6 73.4 74.9
4.1 Dataset and Evaluation Metric
We evaluate our approach on nuScenes Caesar et al. (2020) and Argoverse2 Wilson et al. (2021) datasets.
NuScenes. The nuScenes dataset includes 700, 150, and 150 driving scenes for training, validation, and testing, respectively. The vehicle for data collection is equipped with a 32-beam LiDAR sensor and 6 surrounding-view RGB cameras, thus providing both point clouds and multi-view images. Besides, this dataset annotates more than 1.4 million 3D bounding boxes, across 10 common categories on the street. Both the multi-sensor data and annotation facilitate the exploration of multi-modal 3D object detection. We follow the official policy to mainly evaluate our method on the 3D object detection benchmark in terms of mean average precision (mAP) and nuScenes detection score (NDS). The mAP is averaged over distance thresholds $0.5m$, $1m$, $2m$, and $4m$ on the BEV across $10$ classes. NDS is a weighted average of mAP and other true-positive metrics including mATE, mASE, mAOE, mAVE, and mAAE.
Argoverse2. The Argoverse 2 (AV2) dataset is a large-scale benchmark for perception and prediction in autonomous driving. It comprises 150,000 annotated frames, which is five times larger than the nuScenes dataset, and 1,000 driving scenes. It features two 32-beam LiDARs combined into a 64-beam LiDAR and seven high-resolution surrounding cameras, offering a full 360° field of view and a valid detection range of up to 200 meters, covering an area of 400m × 400m. The dataset is particularly suited for long-range multi-modal object detection tasks. We evaluate our method across 26 categories. For evaluation, in addition to mean Average Precision (mAP), AV2 provides the Composite Detection Score (CDS), a comprehensive metric that combines mAP with other true positive metrics like Average Translation Error (mATE), Average Scale Error (mASE), and Average Orientation Error (mAOE).
@@ 4.2 Experimental Setup
Network Configuration. The image encoder is composed of an image backbone network (i.e., ResNet He et al. (2016) or Swin-Transformer Liu et al. (2021)) and an FPN Lin et al. (2017a) to enrich the multi-scale information. The voxel encoder comprises of a 3D voxel backbone network and a BEV backbone network. For nuScenes, we follow the common practice of using VoxelNet Zhou and Tuzel (2018) as the 3D backbone network, setting the image resolution as $800\times 448$, and setting the voxel size as ($0.075m$, $0.075m$, $0.2m$). The number of object queries is set as $900$. For Argoverse2, there is less literature that can be referred to. Therefore, we follow FSF Li et al. (2024) to use SparseResUNet as the 3D backbone network. The image resolution is $960\times 640$ and the voxel size is ($0.2m$, $0.2m$, $0.2m$). It is worth noting that FSF keeps the output of SparseResUNet without downsampling, while we downsample the output for 4 times to decrease the resolution of BEV features. Since applied for a larger detection range, we use $1600$ object queries on Argoverse2. The configuration of our PoI multi-modal detector on both datasets keeps the same. Specifically, the feature dimension of query embedding is set as $256$. The channel of image and point cloud features is also transformed to $256$ before being fed into the multi-modal decoder. We equally divide the channels of feature maps of each modality into 4 groups, and generate PoIs for each group. This grouping operation improves the capacity of the network Xie et al. (2017) and reduces the computation costs of the dynamic fusion block. Moreover, the first dynamic fusion layer halves the channel of the query embedding, and the second dynamic fusion layer restores the feature dimension. The PoI multi-modal decoder is iteratively applied 6 times, with shared parameters.
Training and Inference. To make fair comparison, we follow the same training pipeline of previous methods Liu et al. (2023b); Cai et al. (2023); Xie et al. (2023); Liang et al. (2022); Yang et al. (2022c); Li et al. (2024), and use the same pretrained models. Specifically, the image encoder is pre-trained on nuImage Caesar et al. (2020) dataset. For nuScenes, the voxel encoder is initialized with the model weights of pretrained TransFusion-L Bai et al. (2022). For Argoverse2, the 3D voxel backbone network is initialized with the model weights of pretrained FSD Fan et al. (2022b), while the BEV backbone network is randomly initialized. For both datasets, our fusion framework is trained for 6 epochs with the AdamW optimizer. For nuScenes, the training sample is resampled by CBGS Zhu et al. (2019) strategy. The initial learning rate is set as 1e-4, adapted with the one-cycle learning rate policy, and the weight decay is set as 0.01. Data augmentation including random flip, random rotation, random translation, random scaling, and random modal masking is adopted. We exploit 8 GPUs for training, with 2 samples on each GPU for nuScenes and 1 sample on each GPU for Argoverse2.
@@
Table 2: Performance comparison with state-of-the-art methods on Argoverse 2 validation split. C-Barrel: construction barrel. MPC-Sign: mobile pedestrian crossing sign. A-Bus: articulated bus. C-Cone: construction cone. V-Trailer: vehicular trailer. Some categories are excluded from the table due to the limited number of instances they contain. However, the average results consider all categories, even those that are omitted. In this experiment, the input voxel size of our PoIFusion is (0.2m, 0.2m, 0.2m), the image resolution is $960\times 640$, and the image backbone is ResNet-50, following the setting of FSF Li et al. (2024).
Methods
Average
Vehicle
Bus
Pedestrian
Box Truck
C-Barrel
Motorcyclist
MPC-Sign
Motorcycle
Bicycle
A-Bus
School Bus
Truck Cab
C-Cone
V-Trailer
Bollard
Sign
Large Vehicle
Stop Sign
Stroller
Bicyclist
mAP Far3D Jiang et al. (2024) 24.4 - - - - - - - - - - - - - - - - - - - -
CenterPoint Yin et al. (2021a) 22.0 67.6 38.9 46.5 40.1 32.2 28.6 27.4 33.4 24.5 8.7 25.8 22.6 29.5 22.4 37.4 6.3 3.9 16.9 0.5 20.1
FSD Fan et al. (2022b) 28.2 68.1 40.9 59.0 38.5 42.6 39.7 26.2 49.0 38.6 20.4 30.5 14.8 41.2 26.9 41.8 11.9 5.9 29.0 13.8 33.4
VoxelNeXt Chen et al. (2023a) 30.5 72.0 39.7 63.2 39.7 64.5 46.0 34.8 44.9 40.7 21.0 27.0 18.4 44.5 22.2 53.7 15.6 7.3 40.1 11.1 34.9
FSF Li et al. (2024) 33.2 70.8 44.1 60.8 40.2 50.9 48.9 28.3 60.9 47.6 22.7 36.1 26.7 51.7 28.1 41.1 12.2 6.8 27.7 25.0 41.6
CMT Yan et al. (2023) 36.1 71.9 41.5 61.2 38.4 62.2 58.2 40.3 55.1 40.3 24.7 47.3 23.6 59.5 23.8 49.6 28.4 6.8 54.7 4.6 50.2
PoIFusion (ours) 40.6 77.6 49.4 70.6 40.7 75.0 67.1 44.0 66.9 55.1 26.9 49.2 32.5 64.6 31.9 59.0 28.1 8.0 44.0 11.4 55.1
CDS Far3D Jiang et al. (2024) 18.1 - - - - - - - - - - - - - - - - - - - -
CenterPoint Yin et al. (2021a) 17.6 57.2 32.0 35.7 31.0 25.6 22.2 19.1 28.2 19.6 6.8 22.5 17.4 22.4 17.2 28.9 4.8 3.0 13.2 0.4 16.7
FSD Fan et al. (2022b) 22.7 57.7 34.2 47.5 31.7 34.4 32.3 18.0 41.4 32.0 15.9 26.1 11.0 30.7 20.5 30.9 9.5 4.4 23.4 11.5 28.0
VoxelNeXt Chen et al. (2023a) 23.0 57.7 30.3 45.5 31.6 50.5 33.8 25.1 34.3 30.5 15.5 22.2 13.6 32.5 15.1 38.4 11.8 5.2 30.0 8.9 25.7
FSF Li et al. (2024) 25.5 59.6 35.6 48.5 32.1 40.1 35.9 19.1 48.9 37.2 17.2 29.5 19.6 37.3 21.0 29.9 9.2 4.9 21.8 18.5 32.0
CMT Yan et al. (2023) 27.8 62.2 33.6 46.8 30.8 47.3 47.6 30.2 43.1 29.8 18.9 38.4 16.9 42.5 17.1 34.5 21.1 5.0 43.0 3.2 40.4
PoIFusion (ours) 31.6 66.5 40.8 54.8 33.0 58.4 54.6 28.7 55.0 42.8 20.4 40.0 24.7 47.8 23.5 42.3 21.6 5.8 35.0 8.7 42.8
4.3 Comparison with State-of-the-art Methods
4.3.1 Results on nuScenes
@@
In Table 1, we make a comprehensive comparison on nuScenes 3D object detection benchmark. The compared methods are divided into three groups: camera-based methods (“C”), LiDAR-based methods(“L”), and multi-modal ones(“C+L”). All the compared multi-modal 3D object detection approaches are without temporal information. Both the test time augmentation and the model ensembling techniques are not used in this comparison. Broadly speaking, the multi-modal methods outperform the methods that leverage only one single modality, validating the benefits of integrating the semantic-intensive RGB image and location-aware LiDAR point clouds.
With the same image resolution and the same voxel size, our proposed PoIFusion achieves state-of-the-art performance on the nuScenes dataset. Specifically, equipped with ResNet-50, our method achieves 73.2% NDS / 71.2% mAP on the nuScenes validation set. Upgrading the image backbone to a more powerful Swin-T Liu et al. (2021), the performance improves to 73.6% NDS and 71.7% mAP. We also submit the inference results to the official test server, and it reports that our PoIFusion with Swin-T image backbone achieves 74.9% NDS and 73.4% mAP, which makes an absolute improvement of 0.8% NDS and 1.4% mAP over the previous best results Yan et al. (2023).
Compared to unified-view approaches Liang et al. (2022); Liu et al. (2023b); Li et al. (2022a), our PoIFusion method preserves modal-specific information more effectively, leading to a significant improvement over the representative BEVFusion model Liu et al. (2023b), with gains of 2.0% in NDS and 3.2% in mAP on the test set. Additionally, in contrast to recent work such as ObjectFusion Cai et al. (2023), which generates object-centric features by fusing regional features through RoI Pooling, our method leverages multi-modal fusion at adaptive Points of Interest (PoIs), resulting in an improvement of 1.6% in NDS and 2.4% in mAP. The use of PoIs enhances the flexibility of sampling locations and enables fine-grained feature fusion, contributing to the superior performance of our approach.
4.3.2 Results on Argoverse2
To study the performance of long-range 3D object detection, which is of significance to ensure safe driving, we further validate our approach on Argoverse2 dataset. The performance comparison is presented in Table 2. Compared to nuScenes, the long-tail category distribution of Argoverse2 emphasizes the importance of semantic-sensitive image features, while the precise localization of such a large detection range heavily relies on LiDAR sensor. Such facts make multi-modal fusion a promising choice for better 3D object detection. As shown in the table, our method consistently outperforms other competitors to deal with the challenge of long-range 3D object detection. Remarkably, the proposed PoIFusion sets a state-of-the-art record on this benchmark, achieving 40.6% mAP and 31.6% CDS. Note that the performance gap between our method and the strongest competitor CMT Yan et al. (2023) on Argoverse2 is more significant than that of nuScenes, further demonstrating the advantage of our PoI-based fusion to extract object-relevant features for addressing more challenge detection tasks.